Mr maneno is considering investing in a risky project which would be added to an existing portfolio. He foresees five possible status of the economy as follows.
Status of the probability return on return on proposed
Economy of existing portfolio investment
A 0.2 16% 12 %
B 0.4 18% 11 %
C 0.2 20% 10 %
D 0.1 22% 9%
E 0.1 24% 8%

The risk free rate of interest is 9%p.a.
Required:
a) Determine whether the proposed project is acceptable. (10 Marks)
b) Explain to Mr Maneno on the usefulness of CAPM in project appraisal.
( 5 Marks)
1
Expert's answer
2016-08-16T08:42:03-0400
a) kj = kRF + (kM – kRF )*bj, where kj is the required rate of return on stock j; kRF is the risk-free rate of return; kM is the required rate of return on the market portfolio, bj is the beta coefficient of stock.
Economy A = 9 + (16 - 9)*0.2 = 10.4%, not accepted.
Economy B = 9 + (18 - 9)*0.4 = 12.6%, accepted.

Economy C = 9 + (20 - 9)*0.2 = 11.2%, accepted.
Economy D = 9 + (22 - 9)*0.1 = 10.3%, accepted.
Economy E = 9 + (24 - 9)*0.1 = 10.5%, accepted.
Overally the project is acceptable.

b) The usefulness of CAPM in project appraisal is explained by the next reasons:
1) It considers only systematic risk, reflecting a reality in which most investors havediversified portfolios from which unsystematic risk has been essentially eliminated.
2) It generates a theoretically-derived relationship between required return and systematicrisk which has been subject to frequent empirical research and testing.
3) It is generally seen as a much better method of calculating the cost of equity than thedividend growth model (DGM) in that it explicitly takes into account a company’s levelof systematic risk relative to the stock market as a whole.
4) It is clearly superior to the WACC in providing discount rates for use in investmentappraisal.